OpenStudy (anonymous):

Solve the equation x/2 = 4 4 2 8 0

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

\[\frac{x}{2} = 4\]\[x = 2(4)\]

OpenStudy (accessdenied):

So, basically, we want to take this red 2: \( \large \frac{x}{\color{red}{2}} = 4 \) and "get rid" of it. To do this using Math, however, we cannot just remove something without breaking equality. So, we must multiply by two to both sides to cancel with the division, thus leaving x on the LHS and leaving 4*2, or 8 \( \large{ 2 * \frac{x}{2} = 2 * 4 \\ x = 2*4 \\ x = 8 } \)